[Brahms-l] Shift report 20050605 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 05 2005 - 23:55:22 EDT
Shifter leader was HHD
Shifters were 


1. A very nice store was produced and kept from 1800 and onwards. However to begin with, RHIC had to recover from a cooling failure.

2. Data runs with FS@2.3A1723, MRS@60A700: 15240-15249
   'Zero-field' run with D1 at A861 and D2-D4 at 0: 15250
   No pedestals taken.

3. No detector problems.
   

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 5 23:55:50 2005

This archive was generated by hypermail 2.1.8 : Sun Jun 05 2005 - 23:56:24 EDT